# OpenStereo: A Comprehensive Benchmark for Stereo Matching and Strong Baseline

### v6

@@ 4.2 Denoising Stereo Matching
MobilenetV2 100 [42] CNN 70.58G 2.78M 0.7737
MobilenetV2 100 [42] CNN 70.58G 2.78M 0.6155
MobilenetV2 100 [42] CNN 70.58G 2.78M 0.6155
MobilenetV2 120d [42] CNN 85.93G 5.21M 0.5573
EfficientNetV2 [48] CNN 157.52G 24.92M 0.5207
RepViT [50] Trans. 101.35G 5.64M 0.5858
MPViT [27] CNN&Trans. 283.35G 13.33M 0.5113
4.2.1 LR_scheduler and Data Augmentation
As shown in Table 1, MultiStepLR yields an EPE of 0.6839, while OneCycleLR achieves a lower EPE of 0.6155. This substantial difference underscores the crucial role of selecting an appropriate learning rate scheduler for stereo matching. The superior performance of OneCycleLR indicates its potential to improve model accuracy and robustness, making it a preferable choice over MultiStepLR for training stereo-matching models. Although five data augmentation techniques—random crop, color augmentation, eraser transform, flip, and spatial transform—are commonly used in stereo matching [33, 60], their empirical efficacy specifically for stereo matching has not been thoroughly explored. This study investigated these data augmentation strategies to address this gap. Visualizations of these augmentations are provided in Figure 4 for further insight. As shown in Table 1, most data augmentation strategies, except for the combined use of HSFlip and random crop, lead to a decline in the model’s EPE metric of SceneFlow. This is because stereo matching involves pixel-level matching, and these data augmentations (color augmentation and spatial transform) can affect the alignment of pixels. The combination of ColorAug, Erase, and Scale (CES) shows the best generalization performance on KITTI2015, with the lowest EPE of 1.56 and D1_all of 7.64, although it increases the EPE on SceneFlow to 0.7240. These findings underscore the importance of selecting appropriate data augmentation methods to enhance model accuracy and robustness.
@@ 4.2.2 Feature Extraction
Table 3: Ablation study of cost volume on the finalpass of SceneFlow test datasets [35]. Gwc represents Group-wise correlation volume [18]. Cat stands for Concatenation volume [6]. G8-C16, G16-C24, and G32-C48 combine Gwc volume and Cat volume [18]. Channel and Dims represent the channel and dimensions of the cost volume, respectively. FLOPs are calculated at a resolution of $544\times 960544 × 960$.
@@ 4.2.3 Cost Construction
Table 4: Ablation study of disparity regression and refinement on the finalpass of SceneFlow test datasets [35]. ArgMin refers to Differentiable ArMin. Context stands for ContextUpasmple [33, 60].
@@ 4.2.4 Disparity Regression and Refinement
The Differentiable ArgMin [25, 6, 18, 47, 59, 15, 70, 46, 52] introduced by GCNet [25], calculates initial disparity by converting matching costs into probabilities via softmax and then computing a weighted sum of these probabilities across all disparity levels. As shown in Table 4, various strategies show differing impacts on model performance in this ablation study on disparity refinement for the SceneFlow test datasets. Without refinement, the model has an EPE of 0.76. RGBRefine and Context methods slightly improve EPE to 0.72 and 0.71, respectively, with a modest increase in computational resources. Combining these methods further reduces EPE to 0.69, indicating marginal benefits from their integration. However, employing ConvGRU refinement substantially improves EPE 0.46, albeit at a significant cost in computational complexity (3023.88 Gflops) and model size (12.51M). This highlights a trade-off between accuracy improvements and increased computational demands.
4.3 Analysis and Discussion
4.3.1 Necessity of Comprehensive Ablation Study
In the evolving landscape of deep stereo matching, comprehensive ablation studies play a pivotal role in deciphering the effectiveness of different components and strategies. A thorough ablation study goes beyond mere performance metrics; it uncovers the underlying mechanics of different algorithms, revealing their strengths and weaknesses in various scenarios. For instance, different data augmentation techniques may yield contrasting effects on the model’s ability to match stereo images accurately. Similarly, the impact of varying backbones, cost volume configurations, and disparity regression methods on the overall performance can be profound. Understanding the specific contributions of each component is crucial for building more efficient and effective stereo-matching systems.
4.3.2 Necessity of A Strong Baseline
A strong baseline in deep stereo-matching research is critical for several key reasons. First, it serves as a vital reference point, enabling a clear assessment of new methods against an established standard. Second, a strong baseline allows for precise evaluation of the impact of specific changes, whether they are new data augmentation methods, different network architectures, or innovative disparity estimation techniques. This helps in isolating and understanding the contribution of each component to the overall performance. Additionally, a solid baseline ensures fair and meaningful comparisons across studies, providing a common ground for evaluating different research outcomes. This is crucial for maintaining consistency and validity in comparative analyses. In summary, a strong baseline is essential for meaningful advancements in deep stereo matching, ensuring that new developments are substantial, accurately assessed, and broadly applicable.
@@ 5.1 Pipeline
Table 5: Results on SceneFlow [35], KITTI 2012 [14], KITTI 2015 [36] leaderboard, and DrivingStereo [66]. All results on DrivingStereo [66] are derived using the OpenStereo. Underline refers to evaluation in the non-occluded regions only STTR [31]. Bold: Best.
Method SceneFlow KITTI 2012 KITTI 2015 DrivingStereo
EPE 3-noc 3-all D1-bg D1-fg D1-all EPE D1-all
STTR [31] 0.43 - - 1.70 3.61 2.01 OOM OOM
PSMNet [6] 1.09 1.49 1.89 1.86 4.62 2.32 1.19 2.26
GwcNet [18] 0.76 1.32 1.70 1.74 3.93 2.11 0.99 1.36
CFNet [47] 1.04 1.23 1.58 1.54 3.56 1.88 0.98 1.46
AANet [61] 0.87 1.91 2.42 1.65 3.96 2.03 2.91 15.16
Mobilestereo-3D [46] 0.80 - - 1.75 3.87 2.10 1.06 1.61
COEX  [2] 0.68 1.55 1.93 1.74 3.41 2.02 1.34 2.70
FADNet++  [52] 0.76 - - 1.99 3.18 2.19 1.44 5.15
CascadeStereo [15] 0.72 - - 1.59 4.03 2.00 1.31 2.84
IGEV-Stereo [60] 0.47 1.12 1.44 1.38 2.67 1.59 1.06 1.50
GANet+ADL [63] 0.50 0.98 1.29 1.38 2.38 1.55 - -
NMRF-Stereo [17] 0.45 1.01 1.35 1.28 3.13 1.59 - -
Selective-IGEV [53] 0.44 1.07 1.38 1.33 2.61 1.55 - -
MoCha-Stereo [8] 0.41 1.06 1.35 1.36 2.43 1.53 - -
StereoBase(Ours) 0.34 1.00 1.26 1.28 2.26 1.44 1.15 2.19
@@ 5.2 Comparison with State-of-the-art Methods
In our comprehensive evaluation, we benchmarked StereoBase against current state-of-the-art methods on SceneFlow [35], KITTI2012 [14], 2015 [36], and DrivingStereo [66]. On the SceneFlow [35] test set, we achieve a new SOTA EPE of 0.34. The quantitative comparisons, as summarized in Table 5, clearly illustrate the edge of StereoBase in handling complex stereo-matching scenarios with greater precision. Further, we submitted our results to the KITTI2012 [14] and 2015 [36] leaderboards, where StereoBase outperformed all published methods across all metrics. On KITTI2015 [36], our StereoBase outperforms IGEV [60] by 9.43% on D1-all metric, respectively. In addition, we evaluate the generalization performance of StereoBase. As shown in Table 6, StereoBase exhibited exceptional performance in a zero-shot setting. This evaluation further validates the adaptability and potential of StereoBase in handling diverse and challenging stereo vision tasks.
Table 6: Cross-domain evaluation on Middlebury, ETH3D, and KITTI all training sets. All methods are only trained on the Scene Flow datatest. Middlebury is tested on half-resolution. The model with indicates the implementation of OpenStereo. Bold: Best.
Method KITTI2012 KITTI2015 Middlebury ETH3D
D1-all(%) EPE D1-all(%) EPE bad 2.0(%) bad 1.0(%)
STTR 49.72 6.80 40.26 6.16 OOM 38.89
PSMNet 30.51 4.68 32.15 5.99 33.53 18.02
CFNet 13.64 2.27 12.09 2.89 23.91 7.67
AANet 7.23 1.27 7.72 1.41 22.45 18.77
Mobilestereo-2D 18.34 2.45 21.21 2.78 34.04 13.89
Mobilestereo-3D 18.96 2.79 19.69 3.40 29.32 13.71
GwcNet 23.05 2.76 25.19 3.58 29.87 14.54
COEX 12.08 1.80 11.00 2.48 25.17 11.43
FADNet++ 11.31 1.77 13.23 2.97 24.17 25.53
CascadeStereo 11.83 1.83 12.03 2.69 27.27 11.68
IGEV 4.88 0.98 5.16 1.18 8.47 3.53
StereoBase 4.85 0.99 5.35 1.18 9.76 3.12
Figure 6: Visualization results on KITTI2015 dataset.
@@ 6 Conclusion
Acknowledgements. This work was supported by the National Natural Science Foundation of China under Grant 62373356 and the Open Projects Program of State Key Laboratory of Multimodal Artificial Intelligence Systems.

### v7

@@ 4.4.1 LR_scheduler and Data Augmentation
MobilenetV2 [37] CNN 70.58G 2.78M 0.77
MobilenetV2 [37] CNN 70.58G 2.78M 0.62
MobilenetV2 [37] CNN 70.58G 2.78M 0.62
MobilenetV2 [37] CNN 85.93G 5.21M 0.56
EfficientNetV2 [44] CNN 157.52G 24.92M 0.52
RepViT [46] Trans. 101.35G 5.64M 0.59
MPViT [24] CNN&Trans. 283.35G 13.33M 0.51
Table 2: Ablation study on Backbones Selection. MobilenetV2 refers to MobilenetV2 120d. Bold: Best.
@@ 4.4.2 Feature Extraction
Table 3: Ablation study on Cost Construction. Gwc represents Group-wise correlation volume [15]. Cat stands for Concatenation volume [5]. G8-C16, G16-C24, and G32-C48 combine Gwc volume and Cat volume [15]. Channel and Dims represent the channel and dimensions of the cost volume, respectively. Bold: Best.
@@ 4.4.3 Cost Construction
Table 4: Ablation study on Disparity Regression and Refinement. ArgMin refers to Differentiable ArMin. Context stands for ContextUpasmple [29, 55]. Bold: Best.
@@ 4.4.4 Disparity Regression and Refinement
The Differentiable ArgMin [21, 5, 15, 42, 54, 13, 64, 41, 48] introduced by GCNet [21], calculates initial disparity by converting matching costs into probabilities via softmax and then computing a weighted sum of these probabilities across all disparity levels. As shown in Table 4, various strategies show differing impacts on model performance in this ablation study on disparity refinement for the SceneFlow test datasets. Without refinement, the model has an EPE of 0.76. RGBRefine and Context methods slightly improve EPE to 0.72 and 0.71, respectively, with a modest increase in computational resources. Combining these methods further reduces EPE to 0.69, indicating marginal benefits from their integration. However, ConvGRU refinement substantially improves EPE to 0.46, albeit at a significant cost in computational complexity (3023.88 Gflops) and model size (12.51M). This highlights a trade-off between accuracy improvements and increased computational demands.
@@ 5 A Strong Pipeline: StereoBase
A strong baseline in deep stereo-matching research is critical for several key reasons. First, it serves as a vital reference point, enabling a clear assessment of new methods against an established standard. Second, a strong baseline allows for precise evaluation of the impact of specific changes, whether they are new data augmentation methods, different network architectures, or innovative disparity estimation techniques. This helps in isolating and understanding the contribution of each component to the overall performance. Additionally, a solid baseline ensures fair and meaningful comparisons across studies, providing a common ground for evaluating different research outcomes. This is crucial for maintaining consistency and validity in comparative analyses.
Method SceneFlow KITTI 2012 KITTI 2015
EPE 3-noc 3-all D1-bg D1-fg D1-all
PSMNet [5] 1.09 1.49 1.89 1.86 4.62 2.32
GwcNet [15] 0.76 1.32 1.70 1.74 3.93 2.11
CFNet [42] 1.04 1.23 1.58 1.54 3.56 1.88
CascadeStereo [13] 0.72 - - 1.59 4.03 2.00
IGEV-Stereo [55] 0.47 1.12 1.44 1.38 2.67 1.59
GANet+ADL [57] 0.50 0.98 1.29 1.38 2.38 1.55
NMRF-Stereo [14] 0.45 1.01 1.35 1.28 3.13 1.59
Selective-IGEV [49] 0.44 1.07 1.38 1.33 2.61 1.55
MoCha-Stereo [7] 0.41 1.06 1.35 1.36 2.43 1.53
StereoBase (Ours) 0.34 1.00 1.26 1.28 2.26 1.44
Table 5: Results on SceneFlow [30], KITTI 2012 [12], and KITTI 2015 [31] leaderboard. Bold: Best.
Figure 6: Visualization results on KITTI2015 dataset.
@@ 5.1 Pipeline
Method K12 K15 Midd ETH3D
STTR [28] 49.72 40.26 OOM 38.89
PSMNet [5] 30.51 32.15 33.53 18.02
CFNet [42] 13.64 12.09 23.91 7.67
AANet [56] 7.23 7.72 22.45 18.77
M2D [41] 18.34 21.21 34.04 13.89
M3D [41] 25.14 22.44 29.32 13.71
GwcNet [15] 23.05 25.19 29.87 14.54
COEX [2] 12.08 11.01 25.17 11.43
FADNet++ [48] 11.31 13.23 24.07 22.48
CStereo [13] 11.86 12.06 27.39 11.62
IGEV [55] 4.88 5.16 8.47 3.53
StereoBase (Ours) 4.85 5.35 9.76 3.12
Table 6: Cross-domain evaluation on Middlebury, ETH3D, and KITTI all training sets. All methods are only trained on the Scene Flow dataset. Middlebury is tested on half-resolution. The model with indicates the implementation of OpenStereo. Bold: Best.
@@ 5.2 Comparison with State-of-the-art Methods
Implementation details Training is driven by the smooth L1 [5] loss function. We utilize the AdamW optimizer in conjunction with the OneCycle learning rate scheduler. For SceneFlow [30], the total training epochs is set to 90. On the KITTI dataset, we finetune the pre-trained model using a combined dataset of KITTI2012 [12] and 2015 [31] training image pairs. The total epoch is 500. (More implementation details in the Supplementary Material).
In our comprehensive evaluation, we benchmarked StereoBase against current state-of-the-art methods on SceneFlow [30], KITTI2012 [12], and KITTI2015 [31]. On the SceneFlow [30] test set, we achieve a new SOTA EPE of 0.34. The quantitative comparisons, as summarized in Table 5, clearly illustrate the edge of StereoBase in handling complex stereo-matching scenarios with greater precision. Further, we submitted our results to the KITTI2012 [12] and 2015 [31] leaderboards, where StereoBase outperformed all published methods across all metrics. On KITTI2015 [31], our StereoBase outperforms IGEV [55] by 9.43% on D1-all metric, respectively. In addition, we evaluate the generalization performance of StereoBase. As shown in Table 6, StereoBase exhibited exceptional performance in a zero-shot setting. This evaluation further validates the adaptability and potential of StereoBase in handling diverse and challenging stereo vision tasks.